# A deconstruction of methods to derive one-point lensing statistics

### v1

@@ 3.5 Maps and power spectrum
Figure 11: Shear maps for the angular resolution of 1.76 arcsec.
@@ 4 Conclusions
In this paper, we investigated the effectiveness of the approximate methods PINOCCHIO and turboGL in accurately reproducing lensing probability density functions (PDFs). Our simulations assume a universe composed solely of dark matter, excluding baryonic effects. We studied the impact of gravitational lensing in both weak and strong regimes, differentiating the contributions from the three types of images. We thoroughly compared the results of these approximate methods with those obtained from $N$-body simulations. Key findings include:
@@
    For supernovae and standard sirens these methods prove useful for statistical analyses, given that strong lensing events are either rare or can be excluded. Within this regime, baryonic effects are expected to have a minor impact.
@@
    These approximate methods struggle to accurately capture the small-scale nonlinear matter field. As the perturbation theory fails to accurately predict small-scale halo clustering and intertwining matter distribution, enhancing the modeling of halo internal structures could improve results. Recent work by Berner et al. (2022) has initiated efforts to reconstruct halo substructure from merger histories, which could be integrated into these methods to add more small-scale power to the simulations. We also observed that non-halo particles minimally impact lensing for $z<3$.
@@
    Our evaluation of the second-to-fourth moments of the lensing PDFs indicated that PINOCCHIO agrees with $N$-body results to within 8–15% in regimes where baryonic effects are minimal.
@@
    The analyses of the moments in Figs. 6 and 7 and the maps (Figs. 911) underscore the need for precise modeling of mass distribution and non-linear clustering of halos with $M>10^{12}M_{\odot}$. The challenges are twofold: ($i$) while there is a moderate consensus on the impact of baryons at the scale of galaxy clusters, their effects in individual galaxies and galaxy groups remain far less understood and more uncertainties (see, e.g., Euclid Collaboration: Castro, T., et al., 2023); ($ii$) lighter objects in filaments require both accurate modeling of their individual profiles and a detailed understanding of their interconnected mass distributions.
@@ Acknowledgements
TC is supported by the Agenzia Spaziale Italiana (ASI) under - Euclid-FASE D Attivita’ scientifica per la missione - Accordo attuativo ASI-INAF n. 2018-23-HH.0, by the PRIN 2022 PNRR project "Space-based cosmology with Euclid: the role of High-Performance Computing" (code no. P202259YAF), by the Italian Research Center on High-Performance Computing Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2, within the activities of Spoke 3, Astrophysics and Cosmos Observations, by the INFN INDARK PD51 grant, and by the FARE MIUR grant ‘ClustersXEuclid’ R165SBKTMA. VM is supported by the Brazilian research agencies FAPES and Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq). MQ is supported by the research agencies Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) project E-26/201.237/2022 and CNPq. We acknowledge the use of the Santos Dumont supercomputer of the National Laboratory of Scientific Computing (LNCC, Brazil) through the project lenssims.
@@
Codes will be shared on reasonable request to the corresponding author. Data products and post-processing routines are available at (Castro et al., 2024).141414github.com/TiagoBsCastro/LensingMethodsDeconstruction
@@ Appendix A PDFs at different angular resolutions
    the convergence peak shifts to lower $|\kappa|$ values and the variance decreases;
@@
    the shear peak shifts to lower $\gamma$ values without a noticeable change in its variance;
@@
    the magnification exhibits a combined effect of i) and ii), where the peak shifts towards $\mu=1$ and the variance decreases.

### v2

@@ 3.5 Maps and power spectrum
Figure 11: Shear maps for the angular resolution of 1.76 arcsec. The differences between PINOCCHIO and $N$-body are visible and similar for both redshifts. PINOCCHIOhalos produces very similar structure to PINOCCHIO  and the differences are mostly on the one-point distribution and thus harder to spot. Compare with Table 4.
@@ 4 Conclusions
In this paper, we investigated the effectiveness of the approximate methods PINOCCHIO and turboGL in accurately reproducing lensing probability density functions. Our simulations assume a universe composed solely of dark matter, excluding baryonic effects. We studied the impact of gravitational lensing in both weak and strong regimes, differentiating the contributions from the three types of images. We thoroughly compared the results of these approximate methods with those obtained from $N$-body simulations. Key findings include:
@@
    For supernovae and standard sirens, these methods prove useful for statistical analyses, given that strong lensing events are either rare or can be excluded. Within this regime, baryonic effects are expected to have a minor impact.
@@
    These approximate methods struggle to capture the small-scale nonlinear matter field accurately. As the perturbation theory fails to accurately predict small-scale halo clustering and intertwining matter distribution, enhancing the modeling of halo internal structures could improve results. Recent work by Berner et al. (2022) has initiated efforts to reconstruct halo substructure from merger histories, which could be integrated into these methods to add more small-scale power to the simulations. We also observed that non-halo particles minimally impact lensing for $z<3$.
@@
    Our evaluation of the second-to-fourth moments of the lensing PDFs indicated that PINOCCHIO agrees with $N$-body results to within 10% in the regime originally proposed by the MeMo method by Quartin et al. 2014, that is, $\Delta m_{\rm cut}\gtrsim-0.4$, suiting it to be used for supernovae lensing. However, it underestimates the second-to-fourth moments by roughly 10, 15, and 20% in a regime where the baryonic effect is yet not dominant, namely $\Delta m_{\rm cut}\gtrsim-1$.
@@
    The analyses of the moments in Figs. 6 and 7 and the maps (Figs. 911) underscore the need for precise modeling of mass distribution and non-linear clustering of halos with $M>10^{12}M_{\odot}$. The challenges are twofold: ($i$) while there is a moderate consensus on the impact of baryons at the scale of galaxy clusters, their effects in individual galaxies and galaxy groups remain far less understood and more uncertain (see, e.g., Castro et al., 2024a); ($ii$) lighter objects in filaments require both accurate modeling of their individual profiles and a detailed understanding of their interconnected mass distributions.
@@ Acknowledgements
It is a pleasure to thank Ben Metcalf for the support with GLAMER software. TC is supported by the Agenzia Spaziale Italiana (ASI) under - Euclid-FASE D Attivita’ scientifica per la missione - Accordo attuativo ASI-INAF n. 2018-23-HH.0, by the PRIN 2022 PNRR project "Space-based cosmology with Euclid: the role of High-Performance Computing" (code no. P202259YAF), by the Italian Research Center on High-Performance Computing Big Data and Quantum Computing (ICSC), project funded by European Union - NextGenerationEU - and National Recovery and Resilience Plan (NRRP) - Mission 4 Component 2, within the activities of Spoke 3, Astrophysics and Cosmos Observations, by the INFN INDARK PD51 grant, and by the FARE MIUR grant ‘ClustersXEuclid’ R165SBKTMA. VM is supported by the Brazilian research agencies FAPES and Conselho Nacional de Desenvolvimento Científico e Tecnológico (CNPq). MQ is supported by the research agencies Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro (FAPERJ) project E-26/201.237/2022 and CNPq. We acknowledge the use of the Santos Dumont supercomputer of the National Laboratory of Scientific Computing (LNCC, Brazil) through the projects lenssims and lensingemu.
@@
Codes will be shared on reasonable request to the corresponding author. Data products and post-processing routines are available at Castro et al. (2024b).161616github.com/TiagoBsCastro/LensingMethodsDeconstruction
@@ Appendix A PDFs at different angular resolutions
    the peak in the convergence PDF shifts to lower $|\kappa|$ values and the variance decreases;
@@
    the peak in the shear PDF shifts to lower $\gamma$ values without a noticeable change in its variance;
@@
    the magnification exhibits a combined effect of i) and ii), where the PDF peak shifts towards $\mu=1$ and the variance decreases.